### v1

# LLM-DetectAIve: a Tool for Fine-Grained Machine-Generated Text Detection

Abstract
[…]
utf8 **footnotetext: Equal contribution.
**footnotetext: Equal contribution.
1 Introduction
The development of advanced large language models (LLMs), such as GPT-4, Claude-3.5, Gemini-1.5, Llama-70b (OpenAI, 2023; Anthropic, 2024; Gemini, 2023; Llama, 2024), significantly increased the prevalence and coherence of machine-generated content. This trend makes it increasingly difficult to differentiate between texts produced by machines and those written by humans  Macko et al. (2023); Wang et al. (2024c, a). As a result, significant concerns are emerging regarding the authenticity and integrity of textual content Crothers et al. (2023); Tang et al. (2024). While many detectors have been developed to address this new challenge Mitchell et al. (2023); Wang et al. (2024b), they often struggle to keep up with the rapid development of LLMs. Generations produced by new models are hard to detect as they become more coherent and represent out-of-distribution instances unseen by detectors during training Macko et al. (2024); Koike et al. (2024). Additionally, the tendency of prompting LLMs to generate more human-like texts or applying LLMs to refine or change the tone of human writings further complicates the detection task.
Figure 1: LLM-DetectAIve Interface: automatic text detection (top) and human detector playground (bottom).
Most prior works focused only on binary detection, determining whether the text is fully generated by a machine or fully written by a human. Such systems might classify machine-polished texts as fully human-written or vice versa. We note that in some scenarios, such as academic writing, it is generally acceptable to use LLMs to enhance human-written text. However, in educational settings, using LLMs to complete entire assignments or even to polish human-written essays is typically prohibited (Susnjak, 2022). Therefore, fine-grained text classification becomes important. For example, detecting the usage of LLMs in text humanization and refinement becomes critical to ensure the fair assessment of students’ genuine knowledge and abilities. Fine-grained identification of MGTs is also important in authorship detection in forensics. To address this problem, we propose a new formulation of the MGT detection task — a multi-way classification with the following labels:
  • I.
    Human-Written: text is created solely by a human author without AI assistance.
  • II.
    Machine-Generated: text is entirely produced by a machine based on input prompts without any human intervention.
  • III.
    Machine-Written Machine-Humanized: text is initially generated by a machine and then subtly modified to appear more human-like. This involves automatically tweaking the MGT to make it more personable, avoid plagiarism, enhance relatability, and increase its overall human quality.
  • IV.
    Human-Written Machine-Polished: text is written by a human author and then is refined or polished by a machine. AI tools are used to correct grammar, improve style, and optimize readability while preserving the original meaning of the text.
Text Class Generator OUTFOX Wikipedia Wikihow Reddit ELI5 arXiv abstract PeerRead
M4GT-Bench
I Human 14,043 14,333 15,999 16,000 15,998 2,847
II davinci-003 3,000 3,000 3,000 3,000 3,000 2,340
gpt-3.5-turbo 3,000 2,995 3,000 3,000 3,000 2,340
cohere 3,000 2,336 3,000 3,000 3,000 2,342
dolly-v2 3,000 2,702 3,000 3,000 3,000 2,344
BLOOMz 3,000 2,999 3,000 2,999 3,000 2,334
gpt4 3,000 3,000 3,000 3,000 3,000 2,344
New Generations
II + III + IV gpt-4o 8,966 8,995 9,000 9,000 9,000 7,527
gemma-7b 8,280 8,985 9,000 9,000 9,000 0
llama3-8b 8,271 8,985 9,000 9,000 9,000 0
llama3-70b 8,577 8,985 9,000 9,000 9,000 0
mixtral-8x7b 17,001 8,985 9,000 9,000 9,000 0
gemma2-9b 0 8,985 9,000 9,000 9,000 0
III gemini1.5 0 1,652 1,601 904 0 0
mistral-7b 0 2,993 3,000 0 0 2,344
IV gemini1.5 0 1,652 1,601 904 2,994 586
mistral-7b 0 2,993 3,000 0 0 2,344
Table 1: Dataset statistical information over four classes across various LLMs. I. Human-Written, II. Machine-Generated, III. Machine-Written Machine-Humanized and IV. Human-Written Machine-Polished. For row II + III + IV, the data is approximately equally distributed across the three classes.
We developed LLM-DetectAIve – a system that accurately distinguishes between different types of text generation and editing. The system aims to uphold academic integrity and ensure a fair evaluation process for both students and researchers. Contributions are as follows:
  • We collected a dataset for training and testing fine-grained detectors.
  • We built several detection models using the collected training data and perform their extensive evaluations.
  • We developed a demo with web interfaces that allow users to input text and detect the fine-grained intervention of LLMs in text generation. It also offers a playground for users to test their abilities of detecting texts with varying degrees of LLM intervention.
2 Dataset
To collect data for a multi-way detector training, we first gather datasets which were curated for binary MGT detection from previous work and then extend the dataset into four labels by introducing new corresponding generations. Sections 2.2 and 2.3 elaborate the prompts used in generation and data cleaning, respectively.
2.1 Data Overview
We build the new dataset based on M4GT-Bench Wang et al. (2024a). It is an MGT evaluation benchmark that encompasses multiple generators and domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays (OUTFOX), and peer reviews (PeerRead). From these sources, we sampled a subset comprising 79,220 human-written texts and 103,075 machine-generated texts. Next, we expand this dataset by (i) collecting additional fully machine-generated texts produced by new LLMs (e.g., GPT-4o), (ii) generating machine-written then machine-humanized data based on fully-MGTs sampled from M4GT-Bench, and (iii) polishing human-written texts by various LLMs for the human-written then machine-polished category. This results in 91,358 fully-MGTs, 103,852 machine-written then machine-humanized texts, and 107,900 human-written then machine-polished texts. Table 1 demonstrates the detailed statistical information of the dataset. For data generation, we used a variety of LLMs, including LLaMA3-8b, LLaMA3-70b Llama (2024), Mixtral 8x7b Jiang et al. (2024), Gemma-7b, Gemma2-9b Team et al. (2024), GPT-4o (OpenAI, 2023), Gemini-1.5-pro (Gemini, 2023), and Mistral-7b Jiang et al. (2023). By incorporating a diverse array of LLMs and domains, we aim to enhance the detection accuracy within actual domains and generators, as well as improve the generalization over unseen inputs.
2.2 Generation Prompts
The concept of the two new text classes encompasses a broad spectrum of potential prompts and methods for generating these texts. We conducted experiments with a variety of prompts to account for possible variations and ideologies embedded within. We primarily used around 5-6 prompts per domain to generate data under “Machine-Written Machine-Humanized” and “Human-Written Machine-Polished” classes. Prompts were chosen uniformly at random for these two new text classes to avoid introducing any artifacts into our dataset. In addition to the “Fully-MGT” class, we used the previous prompts from the M4GT-Bench dataset. For the “Machine-Written Machine-Humanized” class, examples of prompts include Rewrite this text to make it sound more natural and human-written or Rephrase this text to be easy to understand and personable. For the “Human-Written Machine-Polished” class, we used prompts such as Paraphrase the provided text. or Rewrite this text so that it is grammatically correct and flows nicely.. Additionally, we introduced a trailing prompt appended to each randomly selected prompt to prevent undesirable text that the LLM may prepend to its output. An example of a trailing prompt is: “Only output the text in double quotes with no text before or after it. Text: {} Your response:”.
2.3 API Tools & Data Cleaning
For data generation, we utilized multiple APIs, including those from OpenAI, Gemini, Groq, and DeepInfra, to generate a total of 303,110 texts for the three LLM-dependent classes. For each of the two new class generations, we limited text length to 1,500 words to accommodate the context length restrictions of some smaller LLMs and to efficiently manage time and cost. The output of the LLMs occasionally includes formatting such as “Here is the paraphrased text:” and “Sure!” despite instructions in the trailing prompt to exclude any additional output. We removed these phrases in the post-processing with two considerations. On the one hand, this naturally occurs in real-world applications, i.e., humans will remove these irrelevant phrases when they use the target content. Moreover, the presence of these indicative artifacts could significantly impact the detectors’ generalization and the quality of the dataset, given that they are potentially unique for a specific text class.
3 Detection Models
We train three detectors by fine-tuning RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2019) using the dataset collected above. DeBERTa is selected because it was built upon BERT and RoBERTa by incorporating disentangled attention mechanisms and an enhanced mask decoder, improving word representation. To provide instant response to users, we applied DistilBERT. It is a more compact and faster variant of BERT, being 60% faster and 40% smaller than BERT while retaining 97% of BERT’s language understanding capabilities. Table 2 presents the hyperparameters for each model. RoBERTa and DistilBERT were employed in the domain-specific experiments. While due to the inferior performance of DistilBERT to RoBERTa in our preliminary trials, we substituted DistilBERT with DeBERTa in the following experiments (DeBERTa is superior than RoBERTa).
Dataset Detector Learning rate Weight Decay Epochs Batch Size
arXiv RoBERTa 2e-5 0.01 10 16
DistilBERT 2e-5 0.01 10 16
OUTFOX RoBERTa 2e-5 0.01 10 16
DistilBERT 2e-5 0.01 10 16
Full Dataset RoBERTa 5e-5 0.01 10 32
DeBERTa 5e-5 0.01 10 32
Table 2: Detector hyperparameters across models.
4 Experiments
Previous studies have shown that the accuracy of detectors drops substantially for out-of-domain cases (Wang et al., 2024a). To alleviate this, we propose three strategies: (i) train multiple domain-specific detectors, each specifically responsible for detecting inputs from one domain, (ii) train one universal detector using more training data across various domains, and (iii) leverage domain-adversarial neural network (DANN) for domain adaption.
4.1 Domain-specific Detectors
We fine-tuned RoBERTa and DistilBERT using the data of arXiv and OUTFOX across four labels. The ratio of training, validation, and test sets is consistently 70%/15%/15% across all experiments. From the results in Table 3, we see that two detection models demonstrate high accuracy on OUTFOX. Overall, RoBERTa is more robust over diverse domains, accuracy is greater than 95% for both domains, with a small number of mis-classifications occurring between classes with overlapping features, such as Machine-Generated vs. Human-Written, vs. Machine-Polished classes, as the confusion matrices in Figure 2. However, under this design, users are requested to first specify and select the domain of the input text. This will increase the workload of users. To mitigate this, we further train a universal model that does not need the domain router.
4.2 Universal Detector
We fine-tuned RoBERTa and DeBERTa using the full dataset, with the data distribution shown in Table 4. To reduce data imbalance and prevent the detector from favoring any particular class, we excluded some of the original data. The fine-tuning results in Table 5 demonstrate that DeBERTa consistently outperforms RoBERTa across all metrics, indicating superior robustness and reliability. Therefore, we deployed the fine-tuned DeBERTa as the backend detection model for our demo system.
Detector Test Domain Prec Recall F1-macro Acc
RoBERTa arXiv 95.82 95.79 95.79 95.79
OUTFOX 95.67 95.43 95.53 95.65
DistilBERT arXiv 88.98 87.97 87.93 87.79
OUTFOX 96.66 96.65 96.65 96.65
Table 3: Domain-specific performance of RoBERTa and DistilBERT over arXiv and OUTFOX.
Figure 2: Domain-specific confusion matrix of RoBERTa on arXiv (top) and on OUTFOX (bottom).
Domain Human Machine-Generated Machine-Polished Machine-Humanized
Arxiv 15,998 18,000 18,000 18,000
Reddit 16,000 18,904 18,904 18,904
Wikihow 15,999 22,601 22,601 22,601
Wikipedia 14,333 22,615 22,615 22,615
Peerread 2,847 4,684 4,684 4,684
Outfox 14,043 17,000 17,000 17,000
Table 4: Distribution of the data used for fine-tuning universal detectors based on RoBERTa and DeBERTa.
Detector Prec Recall F1-Macro Acc
RoBERTa 94.79 94.63 94.65 94.62
DeBERTa 95.71 95.78 95.72 95.71
Table 5: Detectors performance on the full dataset.
4.3 DANN-based Detector
In domain-specific experiments, we achieved high precision with the domain of the text provided. However, in cross-domain evaluation, the performance is sub-optimal as previous work suggests (Wang et al., 2024a, c). In real-world scenarios, the domain is not always specified. This raises the question: How can we detect the text without knowing its domain? To answer this question, we investigated the use of domain adversarial neural networks Ganin et al. (2016) to train a domain-robust detector. DANN was initially designed to achieve domain adaptation by aligning feature distributions across different domains with three major components:
  • Feature Extractor: to learn a representation of the input data. It can be any feed-forward neural network architecture, such as a CNN for image data.
  • Label Predictor: to predict the class labels based on the extracted features. It is trained using labeled data from the source domain.
  • Domain Classifier: connected to the feature extractor via a gradient reversal layer (GRL), this classifier distinguishes between the source and target domains. It multiplies the gradient by a negative constant during backpropagation, promoting domain-invariant features.
The network is trained using standard back-propagation and stochastic gradient descent, optimizing the label classification loss while intentionally confusing the model regarding the domain by reversing the gradient from the domain classifier. This approach reduces the label classification loss while increasing the domain classification loss. As a result, the Domain-Adversarial Neural Network (DANN) produces a feature vector that is independent of the domain. In this experiment, we fine-tuned the model for four classes across six sources/domains. During training, the DANN acquires both domain- and class-specific knowledge. As demonstrated in Table 6, the application of domain adversarial training to the RoBERTa-based detector enhanced overall performance compared to the fine-tuned RoBERTa discussed in Section 4.2. This finding suggests that decoupling the model from domain-specific features leads to an improvement in its overall performance.
Detector Prec Recall F1-macro Acc
RoBERTa 94.79 94.63 94.65 94.62
DANN+RoBERTa 96.30 95.54 96.06 95.24
Table 6: Performance of domain-specific RoBERTa vs. DANN+RoBERTa. The latter outperforms the former, indicating that decoupling the model from domain-specific features improves the overall performance.
4.4 Comparison with Existing Systems
There are a few MGT detection systems like GPTZero [^2] https://gptzero.me/, ZeroGPT [^3] https://www.zerogpt.com/, Sapling AI detector [^4] https://sapling.ai/ai-content-detector, but none of them supports four-class detection. Only GPTZero can classify mixed text in addition to machine-generated content by adding a third class; however, it limits users to only 40 free scans per day or 10,000 words per month for registered accounts. To compare these systems with LLM-DetectAIve, we evaluated them by randomly sampling 60 machine-generated texts and 60 human texts (10 per source). LLM-DetectAIve achieved a 97.50% accuracy, outperforming GPTZero, ZeroGPT, and Sapling AI, which got 87.50%, 69.17%, and 88.33%, respectively.
5 Demo Web Application
We developed a demo web application with two interfaces: (1) an interface for fine-grained MGT detection; (2) a playground for users.
5.1 Automatic Detection
The automatic detection interface is shown in Figure 1 (top). It allows users to input a text, and then the system responds with the class that the text belongs to. To ensure the prediction accuracy, the length of the submitted text is constrained to 50-500 words since the performance of our detectors drops significantly for shorter texts. Longer texts will be truncated by the tokenizer limited by the context window size of the BERT-like models.
5.2 Human Detector Playground
We developed the human detector playground as an interactive interface to engage users in the classification process. This feature allows users to test their own ability to distinguish between the four text categories. Figure 1 (bottom) shows the playground where users can explore the system, gaining insights into the subtle differences between various types of human and machine-generated texts.
5.3 Deployment and Implementation
The demo web application is deployed using Hugging Face Spaces. It is chosen for its seamless integration with transformer models, ease of use, and robust support for hosting machine learning applications. For implementing the user interface, we use Gradio. The code is publicly available and is licensed under the MIT License.
6 Conclusion and Future Work
In an era of advanced large language models, maintaining the integrity of text presents significant challenges. We presented a system that aims to identify the potential misuse of MGT, accurately differentiating human-written text from various types of machine text. Our system classifies text in a fine-grained manner — human-written, machine-generated, machine-polished, and machine-humanized texts, providing insights into the origins of the text and enhancing the trustworthy applications of LLMs. In future work, we plan to improve the Domain Adversarial Neural Network (DANN) to achieve more robust detection results. Additionally, we will explore the possibility of using a DANN on the text’s generator instead of the text’s domain to generalize detection across different text generators. Using a DANN on the domain and the generator could potentially lead to a truly universal detector. We further aim to add a fifth classification category: machine-written then human-edited text, to enhance our detection capabilities and to provide a more comprehensive analysis of text origins.
Limitations
We acknowledge certain limitations in this work that we plan to address in future research. First, although our work has explored more fine-grained machine-generated text scenarios beyond conventional binary classification, we did not consider a complex scenario where the text is first generated by machine and then manually edited by humans to suit their personal needs. This omission is primarily due to the high cost associated with collecting data that requires human editing. Additionally, issues were identified within the dataset. Specifically, some LLMs associate specific domains with particular formatting styles, such as markdown for lists, bullet points, and headers. This issue was particularly noticeable in the Wikihow and Peerread domains, where the LLMs frequently applied these formatting styles, potentially skewing the data and impacting the accuracy of our classifications. It also remains uncertain whether our system can generalize to detecting models or languages not included in our English only dataset.
Ethical Statement and Broad Impact
Data License
A primary ethical consideration is the data license. We reused pre-existing corpora, such as OUTFOX and Wikipedia, which have been publicly released and approved for research purposes. Furthermore, we generated new data on top of the original data, thereby mitigating concerns regarding data licensing.
Biased and Offensive Language
Considering that our data is generated by large language models, it might contain offensive or biased language.
Positive Impact of Fine-grained Detection
LLM-DetectAIve expands the conventional binary classification in MGT detection to more fine-grained levels, which is more aligned with real-life scenarios. We believe this approach could be applied in various scenarios, such as detecting students’ essays to ensure the originality of their work. Moreover, LLM usage detection may find applications in authorship detection in areas such as forensics.

### v3

@@
 1 Introduction
-The development of advanced large language models (LLMs), such as GPT-4, Claude-3.5, Gemini-1.5, Llama-70b (OpenAI, 2023; Anthropic, 2024; Gemini, 2023; Llama, 2024), significantly increased the prevalence and coherence of machine-generated content.
-This trend makes it increasingly difficult to differentiate between texts produced by machines and those written by humans  Macko et al.
+The development of advanced large language models (LLMs), such as GPT-4, Claude-3.5, Gemini-1.5, Llama-70b (OpenAI, 2023; Anthropic, 2024; Gemini, 2023; Llama, 2024), improved the prevalence and the coherence of machine-generated content.
+This trend makes it increasingly difficult to differentiate between texts produced by machines from such written by humans Macko et al.
 (2023); Wang et al.
-(2024c, a).
-As a result, significant concerns are emerging regarding the authenticity and integrity of textual content Crothers et al.
+(2024b, c).
+As a result, there have been growing concerns about the authenticity and integrity of textual content Crothers et al.
 (2023); Tang et al.
@@
 (2023); Wang et al.
-(2024b), they often struggle to keep up with the rapid development of LLMs.
-Generations produced by new models are hard to detect as they become more coherent and represent out-of-distribution instances unseen by detectors during training Macko et al.
+(2024a), they often struggle to keep up with the rapid development of LLMs.
+Generations produced by new models are hard to detect as they become more coherent and represent out-of-distribution instances, compared to what detecting systems saw during training Macko et al.
 (2024); Koike et al.
 (2024).
-Additionally, the tendency of prompting LLMs to generate more human-like texts or applying LLMs to refine or change the tone of human writings further complicates the detection task.
+Moreover, the use of prompting to generate more human-like texts or applying LLMs to refine or change the tone of human writings further complicates detection.
 Figure 1: LLM-DetectAIve interface: automatic text detection (top) and human detector playground (bottom).
-Most prior works focused only on binary detection, determining whether the text is fully generated by a machine or fully written by a human.
-Such systems might classify machine-polished texts as fully human-written or vice versa.
-We note that in some scenarios, such as academic writing, it is generally acceptable to use LLMs to enhance human-written text.
-However, in educational settings, using LLMs to complete entire assignments or even to polish human-written essays is typically prohibited (Susnjak, 2022).
-Therefore, fine-grained text classification becomes important.
-For example, detecting the usage of LLMs in text humanization and refinement becomes critical to ensure the fair assessment of students’ genuine knowledge and abilities.
-Fine-grained identification of MGTs is also important in authorship detection in forensics.
-To address this problem, we propose a new formulation of the MGT detection task — a multi-way classification with the following labels:
+Most prior work on detecting machine-generated text focused on binary detection, i.e., predicting whether the text is generated by a machine or written by a human.
+This dichotomy leaves no space for mixed categories of human-machine collaboration.
+However, we argue for the need for additional categories, as machine-polishing of human-written text is acceptable in certain cases (e.g., for academic papers), but not in other (e.g., in education).
+In education, using LLMs to complete entire assignments or even to polish human-written essays is typically prohibited (Susnjak, 2022).
+Therefore, it is important to perform fine-grained text classification.
+For example, detecting the use of LLMs in text humanization and refinement becomes critical to ensure the fair assessment of students’ genuine knowledge and abilities.
+Fine-grained human/machine identification is also important for authorship detection in digital forensics.
+A contemporary work by Zhang et al.
+(2024) also highlighted the importance of identifying LLM contributions as coauthors in a human-AI collaborative text.
+To address this problem, we propose a new formulation of problem, as multi-way classification with the following labels:
   • I.
-    Human-Written: text is created solely by a human author without AI assistance.
+    Human-Written: the ext is created solely by a human author without GenAI assistance.
   • II.
-    Machine-Generated: text is entirely produced by a machine based on input prompts without any human intervention.
+    Machine-Generated: the text is entirely produced by a machine based on input prompts without any human intervention.
   • III.
-    Machine-Written Machine-Humanized: text is initially generated by a machine and then subtly modified to appear more human-like.
-This involves automatically tweaking the MGT to make it more personable, avoid plagiarism, enhance relatability, and increase its overall human quality.
+    Machine-Written Machine-Humanized: the text is initially generated by a machine and then subtly modified to appear more human-like.
+This involves automatically tweaking the LLM to make the output appear more human.
   • IV.
-    Human-Written Machine-Polished: text is written by a human author and then is refined or polished by a machine.
-AI tools are used to correct grammar, improve style, and optimize readability while preserving the original meaning of the text.
+    Human-Written Machine-Polished: the text is written by a human and then is refined or polished by a machine, e.g., to correct grammar, improve style, and/or optimize readability while trying to preserve the meaning of the original human text.
 Text Class Generator OUTFOX Wikipedia Wikihow Reddit ELI5 arXiv abstract PeerRead
@@
 mistral-7b 0 2,993 3,000 0 0 2,344
-Table 1: Dataset statistical information over four classes across various LLMs.
-I.
+Table 1: Statistics about our datasets across LLMs over the four classes: I.
 Human-Written, II.
@@
 Human-Written Machine-Polished.
-For row II + III + IV, the data is approximately equally distributed across the three classes.
-We developed LLM-DetectAIve – a system that accurately distinguishes between different types of text generation and editing.
-The system aims to uphold academic integrity and ensure a fair evaluation process for both students and researchers.
-Contributions are as follows:
-  • We collected a dataset for training and testing fine-grained detectors.
-  • We built several detection models using the collected training data and perform their extensive evaluations.
-  • We developed a demo with web interfaces that allow users to input text and detect the fine-grained intervention of LLMs in text generation.
-It also offers a playground for users to test their abilities of detecting texts with varying degrees of LLM intervention.
+For row II + III + IV, the data is approximately uniformly distributed across the three classes.
+We further develop LLM-DetectAIve, a system that accurately distinguishes between different types of text generation and editing.
+With this, we aim to uphold academic integrity and ensure a fair evaluation process for both students and researchers.
+Our contributions are as follows:
+  • We reformulate the task as fine-grained multi-way classification.
+  • We collect a dataset for this reformulation using generations from a variety of LLMs.
+  • We build, evaluate, and compare several machine-generated text detectors on our new fine-grained dataset.
+  • We develop a Web-based demo that (i) allows users to input text and to obtain fine-grained classification prediction, and (ii) offers a playground for users to test their ability to detect texts with varying degrees of LLM involvement, according to the above 4-way fine-grained schema.
 2 Dataset
-To collect data for a multi-way detector training, we first gather datasets which were curated for binary MGT detection from previous work and then extend the dataset into four labels by introducing new corresponding generations.
-Sections 2.2 and 2.3 elaborate the prompts used in generation and data cleaning, respectively.
+To collect the dataset for our multi-way fine-grained detector, we first gathered datasets that were curated for binary machine-generated text detection from previous work, and then we extended the data into our four labels by introducing new corresponding generations.
+Sections 2.2 and 2.3 discuss the prompts we used for generation and data cleaning, respectively.
 2.1 Data Overview
-We build the new dataset based on M4GT-Bench Wang et al.
-(2024a).
-It is an MGT evaluation benchmark that encompasses multiple generators and domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays (OUTFOX), and peer reviews (PeerRead).
+We build the new dataset by extending the M4GT-Bench Wang et al.
+(2024b), which is an benchmark dataset for evaluating machine-generation text detectors that encompasses multiple generators and domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays (OUTFOX), and peer reviews (PeerRead).
 From these sources, we sampled a subset comprising 79,220 human-written texts and 103,075 machine-generated texts.
-Next, we expand this dataset by (i) collecting additional fully machine-generated texts produced by new LLMs (e.g., GPT-4o), (ii) generating machine-written then machine-humanized data based on fully-MGTs sampled from M4GT-Bench, and (iii) polishing human-written texts by various LLMs for the human-written then machine-polished category.
-This results in 91,358 fully-MGTs, 103,852 machine-written then machine-humanized texts, and 107,900 human-written then machine-polished texts.
-Table 1 demonstrates the detailed statistical information of the dataset.
+Next, we expanded this dataset by (i) collecting additional machine-generated texts produced by new LLMs (e.g., GPT-4o), (ii) generating machine-written then machine-humanized texts, and (iii) polishing human-written texts using various LLMs.
+This resulted in 91,358 fully-MGTs, 103,852 machine-written then machine-humanized texts, and 107,900 human-written then machine-polished texts.
+Table 1 gives detailed statistics about the dataset.
 For data generation, we used a variety of LLMs, including Llama3-8b, Llama3-70b Llama (2024), Mixtral 8x7b Jiang et al.
-(2024), Gemma-7b, Gemma2-9b Team et al.
-(2024), GPT-4o (OpenAI, 2023), Gemini-1.5-pro (Gemini, 2023), and Mistral-7b Jiang et al.
+(2024), Gemma-7b, Gemma2-9b Team (2024), GPT-4o (OpenAI, 2023), Gemini-1.5-pro (Gemini, 2023), and Mistral-7b Jiang et al.
 (2023).
-By incorporating a diverse array of LLMs and domains, we aim to enhance the detection accuracy within actual domains and generators, as well as improve the generalization over unseen inputs.
+By incorporating a diverse array of LLMs and domains, we aim to enhance the detection accuracy within actual domains and generators, as well as improve generalization.
 2.2 Generation Prompts
-The concept of the two new text classes encompasses a broad spectrum of potential prompts and methods for generating these texts.
-We conducted experiments with a variety of prompts to account for possible variations and ideologies embedded within.
-We primarily used around 5-6 prompts per domain to generate data under “Machine-Written Machine-Humanized” and “Human-Written Machine-Polished” classes.
-Prompts were chosen uniformly at random for these two new text classes to avoid introducing any artifacts into our dataset.
-In addition to the “Fully-MGT” class, we used the previous prompts from the M4GT-Bench dataset.
-For the “Machine-Written Machine-Humanized” class, examples of prompts include Rewrite this text to make it sound more natural and human-written or Rephrase this text to be easy to understand and personable.
-For the “Human-Written Machine-Polished” class, we used prompts such as Paraphrase the provided text.
-or Rewrite this text so that it is grammatically correct and flows nicely..
-Additionally, we introduced a trailing prompt appended to each randomly selected prompt to prevent undesirable text that the LLM may prepend to its output.
-An example of a trailing prompt is: “Only output the text in double quotes with no text before or after it.
+For the Machine-Written Machine-Humanized class, examples of prompts include Rewrite this text to make it sound more natural and human-written or “Rephrase this text to be easy to understand and personable.” For the Human-Written Machine-Polished class, we used prompts such as “Paraphrase the provided text.” or “Rewrite this text so that it is grammatically correct and flows nicely.” Additionally, we introduced a trailing prompt appended to each randomly selected prompt to prevent undesirable text that the LLM may prepend to its output, e.g., “Only output the text in double quotes with no text before or after it.
 Text: {} Your response:”.
+We used 5-6 prompts per domain to generate data for the Machine-Written Machine-Humanized and Human-Written Machine-Polished classes.
+In addition to the Machine-Generated class, we used the original prompts from the M4GT-Bench dataset.
 2.3 API Tools & Data Cleaning
-For data generation, we utilized multiple APIs, including those from OpenAI, Gemini, Groq, and DeepInfra, to generate a total of 303,110 texts for the three LLM-dependent classes.
-For each of the two new class generations, we limited text length to 1,500 words to accommodate the context length restrictions of some smaller LLMs and to efficiently manage time and cost.
-The output of the LLMs occasionally includes formatting such as “Here is the paraphrased text:” and “Sure!” despite instructions in the trailing prompt to exclude any additional output.
+For data generation, we used multiple APIs from OpenAI, Gemini, Groq, and DeepInfra, to generate a total of 303,110 texts for the three LLM-dependent classes.
+For each of the two new class generations, we limited the text length to 1,500 words in order to accommodate the context length restrictions of some smaller LLMs and to efficiently manage time and costs.
+The output of the LLMs occasionally included formatting such as “Here is the paraphrased text:” and “Sure!” despite instructions in the trailing prompt to exclude any additional output.
 We removed these phrases in the post-processing with two considerations.
 On the one hand, this naturally occurs in real-world applications, i.e., humans will remove these irrelevant phrases when they use the target content.
-Moreover, the presence of these indicative artifacts could significantly impact the detectors’ generalization and the quality of the dataset, given that they are potentially unique for a specific text class.
+Moreover, the presence of these indicative artifacts could impact the detectors’ generalization and the quality of the dataset, given that they are potentially unique for a specific text class.
 3 Detection Models
-We train three detectors by fine-tuning RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2019) using the dataset collected above.
-DeBERTa is selected because it was built upon BERT and RoBERTa by incorporating disentangled attention mechanisms and an enhanced mask decoder, improving word representation.
-To provide instant response to users, we applied DistilBERT.
-It is a more compact and faster variant of BERT, being 60% faster and 40% smaller than BERT while retaining 97% of BERT’s language understanding capabilities.
-Table 2 presents the hyperparameters for each model.
-RoBERTa and DistilBERT were employed in the domain-specific experiments.
-While due to the inferior performance of DistilBERT to RoBERTa in our preliminary trials, we substituted DistilBERT with DeBERTa in the following experiments (DeBERTa is superior than RoBERTa).
+We trained three detectors by fine-tuning RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2019).
+DeBERTa is built upon BERT and RoBERTa by incorporating disentangled attention mechanisms and an enhanced mask decoder, which improves word representation.
+Eventually, in the demo, we used DistilBERT, which is a compact and fast variant of BERT: 60% faster and 40% smaller, while retaining 97% of BERT’s language understanding capabilities.
+Table 2 shows the values of the hyper-parameters for each model.
+We used RoBERTa and DistilBERT in our domain-specific experiments.
+However, due to the inferior performance of DistilBERT to RoBERTa in our preliminary trials, we substituted DistilBERT with DeBERTa in the following experiments (DeBERTa is superior to RoBERTa).
 Dataset Detector Learning rate Weight Decay Epochs Batch Size
@@
 DeBERTa 5e-5 0.01 10 32
-Table 2: Detector hyperparameters across models.
-4 Experiments
-Previous studies have shown that the accuracy of detectors drops substantially for out-of-domain cases (Wang et al., 2024a).
+Table 2: Hyper-parameter values across the models.
+4 Experiments and Evaluation
+The previous studies have shown that the accuracy of detectors drops substantially when testing on out-of-domain examples (Wang et al., 2024b).
 To alleviate this, we propose three strategies: (i) train multiple domain-specific detectors, each specifically responsible for detecting inputs from one domain, (ii) train one universal detector using more training data across various domains, and (iii) leverage domain-adversarial neural network (DANN) for domain adaption.
 4.1 Domain-Specific Detectors
-We fine-tuned RoBERTa and DistilBERT using the data of arXiv and OUTFOX across four labels.
-The ratio of training, validation, and test sets is consistently 70%/15%/15% across all experiments.
-From the results in Table 3, we see that two detection models demonstrate high accuracy on OUTFOX.
-Overall, RoBERTa is more robust over diverse domains, accuracy is greater than 95% for both domains, with a small number of mis-classifications occurring between classes with overlapping features, such as Machine-Generated vs.
+We fine-tuned RoBERTa and DistilBERT using the data from arXiv and OUTFOX, using a ratio of training, validation, and test sets of 70%:15%:15%.
+The results are shown in Table 3.
+We can see that both RoBERTa and DistillBERT performed well on OUTFOX.
+Overall, RoBERTa is more robust over diverse domains, with accuracy greater than 95% on both domains, with a small number of mis-classifications occurring between classes with overlapping features, such as Machine-Generated vs.
 Human-Written, vs.
-Machine-Polished classes, as the confusion matrices in Figure 2.
-However, under this design, users are requested to first specify and select the domain of the input text.
-This will increase the workload of users.
-To mitigate this, we further train a universal model that does not need the domain router.
-4.2 Universal Detector
-We fine-tuned RoBERTa and DeBERTa using the full dataset, with the data distribution shown in Table 4.
+Machine-Polished classes, as the confusion matrices in Figure 2 show.
+However, in this setup, the users need to first specify the domain of the input text, which is an extra effort.
+To mitigate this, we further trained a universal model that does not require the user to select the domain.
+4.2 Universal Detectors
+We fine-tuned RoBERTa and DeBERTa using the full dataset; the data distribution for this is shown in Table 4.
 To reduce data imbalance and prevent the detector from favoring any particular class, we excluded some of the original data.
-The fine-tuning results in Table 5 demonstrate that DeBERTa consistently outperforms RoBERTa across all metrics, indicating superior robustness and reliability.
-Therefore, we deployed the fine-tuned DeBERTa as the backend detection model for our demo system.
+The evaluation results in Table 5 indicate that DeBERTa consistently outperforms RoBERTa across all evaluation measures we use.
+Therefore, we deployed the fine-tuned DeBERTa as the backend detection model for our demo.
 Detector Test Domain Prec Recall F1-macro Acc
@@
 OUTFOX 96.66 96.65 96.65 96.65
-Table 3: Domain-specific performance of RoBERTa and DistilBERT over arXiv and OUTFOX.
-Figure 2: Domain-specific confusion matrix of RoBERTa on arXiv (top) and on OUTFOX (bottom).
-Domain Human Machine-Generated Machine-Polished Machine-Humanized
+Table 3: Domain-specific performance for RoBERTa and DistilBERT on arXiv and OUTFOX.
+Figure 2: Domain-specific confusion matrix for RoBERTa on arXiv (top) and on OUTFOX (bottom).
+Machine- Machine- Machine-
+Domain Human Generated Polished Humanized
 arXiv 15,998 18,000 18,000 18,000
@@
 DeBERTa 95.71 95.78 95.72 95.71
-Table 5: Detectors performance on the full dataset.
+Table 5: Detector performance on the full dataset.
 4.3 DANN-Based Detector
-In domain-specific experiments, we achieved high precision with the domain of the text provided.
-However, in cross-domain evaluation, the performance is sub-optimal as previous work suggests (Wang et al., 2024a, c).
-In real-world scenarios, the domain is not always specified.
-This raises the question: How can we detect the text without knowing its domain?
-To answer this question, we investigated the use of domain adversarial neural networks Ganin et al.
+In our domain-specific experiments above, we achieved strong performance when the domain of the text was provided.
+However, in cross-domain evaluation, the performance is sub-optimal as previous work has suggested (Wang et al., 2024b, c).
+In real-world scenarios, the domain would not always be specified, and thus we need a classifier that is as domain-independent as possible..
+Thus, we investigated the use of domain adversarial neural networks Ganin et al.
 (2017) to train a domain-robust detector.
-DANN was initially designed to achieve domain adaptation by aligning feature distributions across different domains with three major components:
-  • Feature Extractor: to learn a representation of the input data. It can be any feed-forward neural network architecture, such as a CNN for image data.
-  • Label Predictor: to predict the class labels based on the extracted features. It is trained using labeled data from the source domain.
-  • Domain Classifier: connected to the feature extractor via a gradient reversal layer (GRL), this classifier distinguishes between the source and target domains.
-It multiplies the gradient by a negative constant during backpropagation, promoting domain-invariant features.
+DANN was initially designed to achieve domain adaptation by aligning representations across different domains with three major components:
+  • Representation Extractor: which builds a representation of the input data; here, we use RoBERTa.
+  • Label Predictor: to predict the class labels based on the representation; it is trained using labeled data from the source domain.
+  • Domain Classifier: connected to the representation via a gradient reversal layer (GRL), it distinguishes between the source and the target domains.
+It multiplies the gradient by a negative constant during back-propagation, promoting domain-invariant representation.
 The network is trained using standard back-propagation and stochastic gradient descent, optimizing the label classification loss while intentionally confusing the model regarding the domain by reversing the gradient from the domain classifier.
-This approach reduces the label classification loss while increasing the domain classification loss.
-As a result, the Domain-Adversarial Neural Network (DANN) produces a feature vector that is independent of the domain.
-In this experiment, we fine-tuned the model for four classes across six sources/domains.
-During training, the DANN acquires both domain- and class-specific knowledge.
-As demonstrated in Table 6, the application of domain adversarial training to the RoBERTa-based detector enhanced overall performance compared to the fine-tuned RoBERTa discussed in Section 4.2.
-This finding suggests that decoupling the model from domain-specific features leads to an improvement in its overall performance.
+This reduces the label classification loss while increasing the domain classification one.
+As a result, the Domain-Adversarial Neural Network (DANN) yields a representation that is independent of the domain.
+In our experiments, we trained the DANN to predict our four classes and to be as confused as possible when predicting the six sources/domains.
+The results are shown in Table 6.
+We can see that using domain adversarial training on top of RoBERTa-enhances the overall performance compared to just fine-tuning RoBERTa as in Section 4.2.
+This suggests that decoupling the model from domain-specific representation leads to an improvement in its overall performance.
 Detector Prec Recall F1-macro Acc
@@
 DANN+RoBERTa 96.30 95.54 96.06 95.24
-Table 6: Performance of domain-specific RoBERTa vs.
+Table 6: Comparing domain-specific RoBERTa vs.
 DANN+RoBERTa.
-The latter outperforms the former, indicating that decoupling the model from domain-specific features improves the overall performance.
-4.4 Comparison with Existing Systems
-There are a few MGT detection systems like GPTZero [^2] https://gptzero.me/, ZeroGPT [^3] https://www.zerogpt.com/, Sapling AI detector [^4] https://sapling.ai/ai-content-detector, but none of them supports four-class detection.
-Only GPTZero can classify mixed text in addition to machine-generated content by adding a third class; however, it limits users to only 40 free scans per day or 10,000 words per month for registered accounts.
-To compare these systems with LLM-DetectAIve, we evaluated them by randomly sampling 60 machine-generated texts and 60 human texts (10 per source).
-LLM-DetectAIve achieved a 97.50% accuracy, outperforming GPTZero, ZeroGPT, and Sapling AI, which got 87.50%, 69.17%, and 88.33%, respectively.
+The latter outperforms the former across all measures, indicating that decoupling the model from domain-specific representation is beneficial.
+4.4 Comparison to Existing Systems
+There are several previously proposed systems for detecting machine-generated text, such as GPTZero, [^2] https://gptzero.me/ ZeroGPT, [^3] https://www.zerogpt.com/ and Sapling AI detector, [^4] https://sapling.ai/ai-content-detector but none of them supports four classes.
+GPTZero is the only one that goes beyond binary classification: it adds a mixed text; however, it limits users to only 40 free runs per day or 10,000 words per month for registered accounts.
+Thus, we could not perform comparison on our entire test dataset.
+Instead, we randomly sampled 60 machine-generated texts and 60 human texts (10 per source) per source.
+In this binary classification setting, LLM-DetectAIve achieved 97.50% accuracy, outperforming GPTZero, ZeroGPT, and Sapling AI, with 87.50%, 69.17%, and 88.33%, respectively.
+4.5 Generalization Evaluation
+To evaluate the generalization ability of our detector on unseen domains and generators, we experimented with testing on two additional datasets: MixSet (Zhang et al., 2024) and IELTS essays written by individuals for whom English is a second language.
+[^5] https://huggingface.co/datasets/chillies/IELTS_essay_human_feedback For the IELTS essays, after deduplication, we randomly sampled 300 (essay problem statement, human-written essay) pairs, and then we produced the corresponding machine-written essays using the problem statements based on Llama3.1-70B.
+We further generated Machine-Written Machine-Humanized and Human-Written Machine-Polished.
+For MixSet, the original dataset contains a total of 3,600 examples, with 300, 300, 600, and 2,400 examples for Human-Written, Machine-Generated, Machine-Written Machine-Humanized and Human-Written Machine-Polished, respectively.
+It involves models such as Llama2-70B and GPT-4, and text covering domains of email content, news, game reviews, and so on.
+The results are shown in Table 7, where we can see that the detector performs much worse on unseen domains and generators, compared to in-domain and in-generator cases.
+The performance on IELTS is better than on MixSet.
+This can be attributed to the inclusion of the OUTFOX data (English native-speaker student essays) in the training data, while the domains and the generators in MixSet are not in the training set.
+The low generalization performance suggests challenges in adapting black-box detectors to the diverse domains and generators in real-world applications.
+Dataset Prec Recall F1-macro Acc
+IELTS 63.74 66.91 66.55 66.91
+MixSet 59.18 64.25 54.95 60.08
+Table 7: Cross-domain evaluation of our detector on unseen domains and generators: IELTS and MixSet.
 5 Demo Web Application
-We developed a demo web application with two interfaces: (1) an interface for fine-grained MGT detection; (2) a playground for users.
+Our demo web application has two interfaces: (i) an interface for fine-grained MGT detection, and (ii) a playground for users.
 5.1 Automatic Detection
@@
 To ensure the prediction accuracy, the length of the submitted text is constrained to 50-500 words since the performance of our detectors drops significantly for shorter texts.
-Longer texts will be truncated by the tokenizer limited by the context window size of the BERT-like models.
+Longer texts will be truncated, as we are limited by the context window size of the BERT-like transformers we use.
 5.2 Human Detector Playground
-We developed the human detector playground as an interactive interface to engage users in the classification process.
-This feature allows users to test their own ability to distinguish between the four text categories.
-Figure 1 (bottom) shows the playground where users can explore the system, gaining insights into the subtle differences between various types of human and machine-generated texts.
+The demo further offers a human detector playground as an interactive interface, which allows users to test their capability to distinguish between the four text categories.
+Figure 1 (bottom) shows a snapshot of the playground interface where the users can try the system, gaining insights into the subtle differences between various types of human-written and machine-generated texts.
 5.3 Deployment and Implementation
-The demo web application is deployed using Hugging Face Spaces.
-It is chosen for its seamless integration with transformer models, ease of use, and robust support for hosting machine learning applications.
-For implementing the user interface, we use Gradio.
-The code is publicly available and is licensed under the MIT License.
+Our demo is deployed on Hugging Face Spaces, which allows seamless integration with transformer models, ease of use, and robust support for hosting machine learning applications.
+For implementing the user interface, we used Gradio.
+The code is publicly available under an MIT license.
 6 Conclusion and Future Work
-In an era of advanced large language models, maintaining the integrity of text presents significant challenges.
-We presented a system that aims to identify the potential misuse of MGT, accurately differentiating human-written text from various types of machine text.
-Our system classifies text in a fine-grained manner — human-written, machine-generated, machine-polished, and machine-humanized texts, providing insights into the origins of the text and enhancing the trustworthy applications of LLMs.
-In future work, we plan to improve the Domain Adversarial Neural Network (DANN) to achieve more robust detection results.
-Additionally, we will explore the possibility of using a DANN on the text’s generator instead of the text’s domain to generalize detection across different text generators.
-Using a DANN on the domain and the generator could potentially lead to a truly universal detector.
-We further aim to add a fifth classification category: machine-written then human-edited text, to enhance our detection capabilities and to provide a more comprehensive analysis of text origins.
+In an era of advanced large language models, maintaining the integrity of text poses significant challenges.
+We presented a system that aims to identify the use of machine-generated text, accurately differentiating human-written text from various types of automatically generated text.
+Unlike previous work, we use a fine-grained classification schema (Human-Written, Machine-Generated, Machine-Written Machine-Humanized, and Human-Written Machine-Polished), which offers insights into the origins of the text, thus enabling trustworthiness.
+In future work, we plan to improve the Domain Adversarial Neural Network (DANN) to improve the results even further.
+We further plan to explore the possibility of using a DANN on the text’s generator instead of the text’s domain to generalize detection across different text generators.
+Using a DANN on both the domain and the generator could potentially lead to a truly universal detector.
+We also aim to expand the classification to include a fifth category: machine-written and human-edited text, enhancing detection capabilities and providing a more comprehensive analysis of text origins.
+To further improve the system, we also plan to address potential biases in the dataset caused by formatting styles linked to specific domains, such as Wikihow and PeerRead, to ensure better robustness across a broader range of human-written content.
+Last but not least, we want to expand the dataset to encompass a diverse set of languages, enabling the development of a robust multilingual detection model.
 Limitations
-We acknowledge certain limitations in this work that we plan to address in future research.
-First, although our work has explored more fine-grained machine-generated text scenarios beyond conventional binary classification, we did not consider a complex scenario where the text is first generated by machine and then manually edited by humans to suit their personal needs.
-This omission is primarily due to the high cost associated with collecting data that requires human editing.
-Additionally, issues were identified within the dataset.
+We acknowledge certain limitations of our work, which we plan to address in future work.
+First, although our work has explored more fine-grained machine-generated text scenarios beyond conventional binary classification, we did not consider a complex scenario where the text is first generated by a machine and then is manually edited by humans to suit their personal needs.
+This is primarily due to the high costs associated with collecting data that requires human editing.
+Moreover, we identified some issues with the dataset.
 Specifically, some LLMs associate specific domains with particular formatting styles, such as markdown for lists, bullet points, and headers.
 This issue was particularly noticeable in the Wikihow and PeerRead domains, where the LLMs frequently applied these formatting styles, potentially skewing the data and impacting the accuracy of our classifications.
-It also remains uncertain whether our system can generalize to detecting models or languages not included in our English only dataset.
+It also remains uncertain whether our system can generalize to detecting models or languages not included in our English-only dataset.
 Ethical Statement and Broad Impact
@@
 We reused pre-existing corpora, such as OUTFOX and Wikipedia, which have been publicly released and approved for research purposes.
-Furthermore, we generated new data on top of the original data, thereby mitigating concerns regarding data licensing.
+Moreover, we generated new data on top of the original data, thereby mitigating concerns regarding data licensing.
 Biased and Offensive Language
-Considering that our data is generated by large language models, it might contain offensive or biased language.
+Considering that our data is generated by large language models, it might contain offensive or biased language; we did not try to control for this, replying on the inetnal safety mechanisms of the LLMs we used.
 Positive Impact of Fine-Grained Detection
-LLM-DetectAIve expands the conventional binary classification in MGT detection to more fine-grained levels, which is more aligned with real-life scenarios.
-We believe this approach could be applied in various scenarios, such as detecting students’ essays to ensure the originality of their work.
-Moreover, LLM usage detection may find applications in authorship detection in areas such as forensics.
+LLM-DetectAIve expands the conventional binary classification in machine-generated text detection to more fine-grained levels, which is more aligned with real-life scenarios.
+We believe this approach could be applied in various scenarios, e.g., for students’ essays to ensure the originality of their work.
+Moreover, LLM usage detection may find applications in authorship detection as well as in digital forensics.